OpenStudy (anonymous):

Need some help please. Obtain the general solution: (D^2+D)y=-cos x

OpenStudy (usukidoll):

is this differential equations?

OpenStudy (anonymous):

Yes

OpenStudy (usukidoll):

what book are you using ... this looks familiar somehow....

OpenStudy (anonymous):

My instructor does not require a text book.. Let me show you how far I've gotten..

OpenStudy (usukidoll):

oh this looks like one of those 1960s problems I had that last semester...they're the only books I know that uses D

OpenStudy (anonymous):

Lol yeah she is a bit old school..

OpenStudy (usukidoll):

is this differential equations I or II? I feel like I should say undetermined coefficients... because of that -cos x thing... so that would be Yp = Acosx +Bsinx but this could also be something else... like variation of parameters...or something

OpenStudy (anonymous):

The full solution is: \[y=C1+C2e^{-x}+\frac{ 1 }{ 2 }\cos x - \frac{ 1 }{ 2 }\sin x\] I can get the first half but the 2nd half is where I need help.

OpenStudy (anonymous):

Yes you're right on.. it is DE I, Linear, non-homogeneous equations by undetermined coeff.

OpenStudy (usukidoll):

ok so we need to find the Yp and that -cosx ... means that's in the form of Yp =Acosx +Bsinx you need to find the first and second derivative of Yp and plug it into the equation... if you have 0 = something on the right we need a multiplier like by x on Yp and do the whole process again.

OpenStudy (usukidoll):

so find the derivatives of the Yp and plug it into the D's D = 1st derivative D^2 = second derivative

OpenStudy (anonymous):

So just to make sure, would the first derivative be: \[y'p=A(-\sin x)+B(\cos x)\] and the Second Der.: \[y''p=A(-\cos x)+B(-\sin x)\]

OpenStudy (usukidoll):

yeah it should... I have to scan something... keep working at this problem

OpenStudy (usukidoll):

oh wait not scan something more like have to ummm do something.. so I'll be back in 15 minutes..so in the meantime plug in those derivatives in D and then we need to find all the cosx and sinx... all the sinx should have a 0 on the right since we don't have a sinx to deal with in the first place. but we do have a cosx and we need to solve to find our A and B.

OpenStudy (usukidoll):

the Y is confusing on the original problem... maybe if we have it as D^2+D = -cosx we can see it easier.

OpenStudy (anonymous):

Ok im just not sure exactly how to plug them in but ill give it a shot..

OpenStudy (usukidoll):

O_O second derivative of Yp goes in D^2 first derivative of Yp goes in D

OpenStudy (usukidoll):

I'll come back.. in 10-15 minutes.... need to wash up

OpenStudy (anonymous):

right, my bad, but then im not sure where to go from there..

OpenStudy (anonymous):

After plugging in, I have: \[[A(-\cos x)+B(-\sin x)]+[A(-\sin x)+B(\cos x)] = - \cos x\]

OpenStudy (anonymous):

I just don't really know how to clean it up lol

OpenStudy (kainui):

First off how do you find D?\[D=\left[\begin{matrix}0 & -1 \\ 1& 0\end{matrix}\right]\] since when you multiply a vector with the basis {sinx, cosx} you will get it's derivative since the derivative is just a linear transformation! Try it now! \[\left(\begin{matrix}3 \\ -2\end{matrix}\right)=3 \sin x -2 \cos x\] So now if you multiply \[\left[\begin{matrix}0 & -1 \\ 1 & 0\end{matrix}\right] \left(\begin{matrix}1 \\ 0\end{matrix}\right)\] this is the same thing as taking the first derivative of sin(x) right?

OpenStudy (anonymous):

Wow I was never shown that way!

OpenStudy (kainui):

Similarly, D^2 will just be the negative identity matrix. No surprise there either?

OpenStudy (kainui):

So now add up D^2+D which is multiplied by your vector y. This equals the other vector, (0,-1) since that is just -cosx. Since it's really easy to find the inverse of a 2x2 matrix, just multiply by the inverse of each to get y!\[\bar y = (D^2+D)^{-1}\left(\begin{matrix}0 \\ -1\end{matrix}\right)\]

OpenStudy (kainui):

I could give more help if you need it, I'm just giving the basic outline here of what's going on.

OpenStudy (anonymous):

Interesting.. it was never explained that way... now im looking for my Yp and once I find it I plug back into the original equation right? after I do that I am confused on how to clean it up or how to find A and B..

OpenStudy (kainui):

Well that would be the slower method of doing it... But from the notation your teacher is using it suggest she wants you to do it the way I'm showing you.

OpenStudy (anonymous):

Ok, can you show me how to do it by your method?

OpenStudy (usukidoll):

WHAT THE HECK HAPPENED HERE?

OpenStudy (usukidoll):

I had old problems like these.... I had to do it the slower method and it worked out so ... if it ain't broke don't fix it.

OpenStudy (anonymous):

Lol this dude tried to help me but then left... Did u see where I plugged them in?

OpenStudy (usukidoll):

\[[A(-\cos x)+B(-\sin x)]+[A(-\sin x)+B(\cos x)] = - \cos x \]

OpenStudy (anonymous):

Yes.. just not sure how to clean up or solve..

OpenStudy (usukidoll):

\[[-A\cos x)+(-B\sin x)]+[(-A\sin x)+B(\cos x)] = - \cos x \]

OpenStudy (usukidoll):

\[-A \cos x + B cosx = -cosx \] \[-Asinx -Bsinx = 0 \]

OpenStudy (usukidoll):

you split them up into two parts.. so factor out the cosx and divide by cosx same goes for sinx

OpenStudy (usukidoll):

\[cosx(-A+B) = -cosx \]

OpenStudy (usukidoll):

\[sinx( -A-B) = 0\]

OpenStudy (anonymous):

Ohhhhhhhhhhhhh.... nice

OpenStudy (usukidoll):

\[-A + B = -1\] \[-A-B=0\]

OpenStudy (usukidoll):

\[-2A = -1 \] \[A = \frac{1}{2}\]

OpenStudy (usukidoll):

\[-\frac{1}{2}-B = 0\]

OpenStudy (usukidoll):

\[-B =\frac{1}{2}\] \[B = \frac{-1}{2}\]

OpenStudy (anonymous):

I see!! Nice thanks.. I guess that was just a brain fart.. I needed to see a couple problems worked out, I've got quite a bit to do and I missed class that day because I was sick.

OpenStudy (usukidoll):

plugging those two values back into the ORIGINAL \[Y_p\] would result in \[\frac{1}{2} \cos x -\frac{1}{2}sinx\]

OpenStudy (usukidoll):

now combine that with the result of \[Y_h\] will become \[y=C1+C2e^{-x}+\frac{ 1 }{ 2 }\cos x - \frac{ 1 }{ 2 }\sin x\]

OpenStudy (anonymous):

You're a lifesaver! Btw, whats your major if you dont mind me asking?

OpenStudy (usukidoll):

MAth :P hence the stupid refine the relation on a strict poset problem being unanswered. I got the solution on paper not sure if it's right but hey at least I tried :P

OpenStudy (usukidoll):

differential equations i is math for mechanical engineers I'm taking another differential equations course next semester that's for Physics majors. ... I think. I love the math, but hate the technical reading.

OpenStudy (anonymous):

Gotcha! Thats awesome! Im actually majoring in Petroleum Eng. Well thanks very much for your help... Hope to see you on here soon cuz I know ill need the help again!